# Mitigating Low-Frequency Bias: Feature Recalibration and Frequency Attention Regularization for Adversarial Robustness

### v1

@@ Abstract
Ensuring the robustness of computer vision models against adversarial attacks is a significant and long-lasting objective. Motivated by adversarial attacks, researchers have devoted considerable efforts to enhancing model robustness by adversarial training (AT). However, we observe that while AT improves the models’ robustness against adversarial perturbations, it fails to improve their ability to effectively extract features across all frequency components. Each frequency component contains distinct types of crucial information: low-frequency features provide fundamental structural insights, while high-frequency features capture intricate details and textures. In particular, AT tends to neglect the reliance on susceptible high-frequency features. This low-frequency bias impedes the model’s ability to effectively leverage the potentially meaningful semantic information present in high-frequency features. This paper proposes a novel module called High-Frequency Feature Disentanglement and Recalibration (HFDR), which separates features into high-frequency and low-frequency components and recalibrates the high-frequency feature to capture latent useful semantics. Additionally, we introduce frequency attention regularization to magnitude the model’s extraction of different frequency features and mitigate low-frequency bias during AT. Extensive experiments showcase the immense potential and superiority of our approach in resisting various white-box attacks, transfer attacks, and showcasing strong generalization capabilities.
@@
DNNs have achieved remarkable success in various applications across diverse domains. However, their susceptibility to inconspicuous adversarial perturbations [1, 2, 3] remains a substantial concern. These subtle perturbations, which are undetectable by the human eye, have the potential threat of resulting in erroneous predictions [4, 5]. As a consequence, the security and robustness of DNNs have become areas of widespread concern within both academic and industrial communities. Adversarial training (AT) [6, 7, 8, 9] is widely acknowledged as one of the most effective techniques to enhance adversarial robustness of models. The AT accomplishes this by strategically integrating meticulously crafted adversarial examples (AEs) into the training process.
@@
Despite the remarkable performance achieved by AT, relying solely on data augmentation using AEs appears insufficient to provide a comprehensive solution. Numerous studies have conducted an exploration of regularization [10, 11] or feature denoising [12, 13] approaches aimed at improving adversarial robustness. Recent advancements based on Fourier analysis [14, 15, 16, 17] have underscored the potential of incorporating frequency analysis during AT to improve robustness. Motivated by these findings, Bu et al. [18] introduced an adaptive module to control frequency preferences, effectively adjusting the low-frequency and high-frequency components of feature representation. Yucel et al. [19] proposed a data augmentation method called HybridAugment that aims to mitigate the dependence of convolutional neural networks (CNNs) on high-frequency components. Furthermore, they propose HybridAugment++ as an attempt to consolidate various frequency-spectrum augmentations. However, it is important to recognize these approaches discard or suppress high-frequency components, which can lead to a significant loss of semantic information in the image [15].
In Figure 1, we illustrate the relationship between the proportion of feature frequency components and prediction accuracy during network inference. A Fourier transform is applied to the feature maps obtained after the initial convolutional layer to explore the impact of varying frequency components on prediction accuracy. By preserving the minimum frequency components across different ratios and subsequently reconstructing the feature maps using the inverse Fourier transform, we can analyze the distinct contributions of various frequency components in feature derivation. We compare WRN32-10 trained with simple training (ST, Fig. 1 left) and adversarial training (AT, Fig. 1 right) on CIFAR-10 dataset. This figure provides empirical evidence that AT significantly improves the robustness of models. Based on the observations from the left in Fig. 1, we can find that adversarial perturbations on images lead to a decline in the model’s prediction accuracy when high-frequency features are present. This suggests that adversarial perturbations primarily affect the high-frequency domain of the features during the network’s inference process. Furthermore, we observe that the robustness achieved by AT does not imply an improved capacity to extract information from high-frequency components. Instead, it tends to bias the model towards relying more on low-frequency information while neglecting vulnerable high-frequency features.
Therefore, we propose a compelling perspective: the robustness of low-frequency and high-frequency features has distinct implications within the inference process of an AT-trained network. Specifically, the robustness of low-frequency features refers to the network’s capacity to effectively capture semantically valuable information for predictions, even when the low-frequency domain is compromised. On the other hand, the robustness of high-frequency features implies a reduced reliance of the network on high-frequency components, potentially resulting in sub-optimal performance in capturing subtle textures and intricate details [20].
@@
To achieve this, we proposed a novel module “High-Frequency Disentanglement and Recalibration (HFDR)”. Firstly, HFDR employs high-pass filters to generate element-wise frequency attention maps, facilitating the disentanglement of features across different frequency domains. Then through the recalibration of vulnerable high-frequency features, the model effectively captures potential semantic cues within the high-frequency domain. The fusion of recalibrated high-frequency features with low-frequency features enhances the network’s ability to extract distinct frequency features. To further mitigate the low-frequency biases induced by AT, we introduce a frequency domain attention regularization. This regularization strategy aims to harmonize the network’s extraction capacity for different frequency domain information. Notably, our proposed model achieves comparable or superior robustness compared to state-of-the-art methods, while incurring minimal additional computational overhead.
@@
Our research underscores a low-frequency bias in conventional adversarial training methods during the network inference process. This robustness of these AT models is primarily established by disregarding vulnerable high-frequency features, which harbor valuable latent semantic information.
We introduce an innovative module called High-Frequency Disentanglement and Recalibration (HFDR). This module enhances the model’s robustness by extracting latent valuable high-frequency cues and mitigates the low-frequency bias issue through incorporating frequency attention mechanisms.
Our model achieves state-of-the-art performance in overall robustness with minimal additional cost and can be integrated with other adversarial training methods to enhance the model’s robustness.
@@ II-A Adversarial Attack
Deep neural networks (DNNs) demonstrate remarkable performance in various tasks by learning intricate relationships between inputs and outputs through complex, non-linear, and high-dimensional mappings. However, DNNs are susceptible to adversarial attacks, in which carefully crafted perturbation can deceive the model, leading to incorrect predictions. Several research studies have been dedicated to investigating the vulnerability of models and have proposed various advanced adversarial attack methods.
@@ II-B Adversarial Training Defense Methods
The objective of the inner maximization is to generate adversarial examples, including attacks such as the PGD attack [21] and other attacks [26, 27, 28, 29, 30]. The outer minimization aims to develop effective training strategies for optimizing network parameters, thereby enhancing the model’s robustness against attacks. Rice et al. [31] conducted a study on the occurrence of overfitting in robust adversarial training and proposed employing a validation set protocol while performing model selection (PGD-AT):
Zhang et al. [32] categorized robust errors into “natural errors” and “boundary errors” to balance robustness and accuracy (TRADES):
Wang et al. [33] explicitly distinguished between misclassified and correctly classified examples during adversarial training (MART):
Wu et al. [34] introduced the Adversarial Weight Perturbation (AWP) to enhance the robustness of the model by considering the change in loss relative to the weight:
Jia et al. [6] introduced a computationally intensive method, Learnable Attack Strategies Adversarial Training (LAS-AT), which involves automatically generating attack from strategy network and strategy set during the training process to enhance robustness:
These methods primarily rely on attack strategy redesign and optimization of the adversarial training process, often overlooking the analysis of feature properties during network inference.
@@ III Methodology
Adversarial training can introduce an inherent low-frequency bias, which impedes the model’s capacity to effectively capture meaningful semantic information across diverse frequency domains. The root cause of this bias lies in the concentration of adversarial perturbations within the high-frequency domain. To tackle this concern, we present a novel “High-Frequency Disentanglement and Recalibration (HFDR)” Module as depicted in Figure 2. The HFDR module is formed of three key components: feature disentanglement, high-frequency feature recalibration, and frequency-based attention regularization.
Throughout the feature disentanglement phase (Section III-A), we generate feature attention maps to selectively focus on distinct frequency domains, facilitating the separation of features into their high-frequency and low-frequency components. The high-frequency feature recalibration stage (Section III-B) recalibrates the high-frequency information to effectively capture valuable semantic features. Additionally, we introduced an innovative regularization technique, termed “Frequency Attention Regularization”, to mitigate the low-frequency bias induced by Adversarial training (Section III-C).
@@ III-A Feature Disentanglement
Feature disentanglement refers to the process of separating and isolating different components or aspects within a given feature representation at a specific layer. In the context of network inference, let represent the feature maps at a pre-defined layer, where , , and are the channel, height, and width dimensions of , respectively. Drawing inspiration from the work of [37], we employ SRM filter [38] to extract high-frequency features. Subsequently, the channels are aligned through a convolution, resulting in the formation of the high-frequency feature representation denoted as . However, recent approaches commonly employ low-pass or high-pass filters for feature disentanglement, resulting in the irreversible loss of information [39, 40].
In order to achieve feature disentanglement while minimizing information loss to the greatest extent possible, we propose employing filters to generate element-wise attention maps for both high-frequency (HF) and low-frequency (LF) semantic sub-features. These attention maps are subsequently utilized to disentangle the features. To avoid the misconception about the enhanced model robustness is attributable to gradient masking, we integrate the Gumbel-Softmax method [41] for generating differentiable attention maps specifically for high-frequency components:
where represents distinct semantic sub-features, and denotes the sigmoid operation. The term incorporates Gumbel noise, defined as with . The term governs the impact of . Additionally, the attention maps for low-frequency features can be computed as .
@@ III-C Frequency Attention Regularization
Information within the feature map is represented at different frequencies: low-frequency conveys the global structure of image features, while high-frequency unveils the local details and structure. However, previous AT methods [31, 34, 33, 18, 46] often induced frequency biases during network inference. Without explicit guidance on network frequency preferences, the network may neglect the separated and refined high-frequency features .
Recall that we obtain the high-frequency attention map for each channel . We propose a straightforward method termed frequency attention regularization (FAR). The FAR aims to enhance the model’s ability to effectively extract both high-frequency and low-frequency features, while simultaneously mitigating inherent low-frequency bias during AT:
where denotes the attention map for the -th channel of the -th image. symbolizes the number of images, and defines the number of channels in the feature maps processed by the HFDR module. Moreover, serves as a balance factor that controls the ratio of information extraction between high-frequency and low-frequency components, while indicates the norm exponent.
@@ III-D Model Training
where controls the influence of , and represents the loss function associated with various AT models. By integrating the HFDR module, we augment the capability of AT models to extract features across diverse frequency domains, thereby improving their adversarial robustness.
@@ IV-A Experimental Setup
Implementation Details: We evaluate the robustness of our approach by conducting experiments on the CIFAR-10 [47], CIFAR-100 [47], and Tiny Imagenet [48] datasets, utilizing the WideResNet34-10 (WRN34-10) [49], AlexNet [17], and ResNet-18 (RN-18) [50] as the baseline models. For the implementation, we insert the HFDR module after the initial convolutional layer. To evaluate the impact of HFDR on improving the robustness of various AT techniques, our model was combined with three AT methods (PGD-AT [31], MART [33], AWP [34]), denoted as, HFDR-AT, HFDR-MART, and HFDR-AWP. We compared our approach with several baseline methods, including, PGD-AT [31], TRADES [32], AWP [34], MART [33], LBGAT [51], LAS-AT [6], FSR [52], FPCM [18], and CFA [53].
We trained WRN34-10 with identical hyperparameters and training details as specified in the original paper [31, 33, 34]. To ensure a fair comparison, we adopted the configuration of PGD-AT [31] as used in the experimental settings of FPCM [18], FSR [52] and CFA [53]. For ResNet18 and AlexNet, we initialized the learning rate to 0.1. The learning rate was subsequently reduced by a factor of 0.1 at the 90th and 95th epochs. The SGD optimizer with a momentum of 0.9 and a weight decay factor of 5e-4 was employed for optimization. In the HFDR hyperparameter settings, we set and for the loss term in Equation 10. The experiments were conducted on two NVIDIA RTX-A4000 GPUs.
@@
| HFDR-AT | Ours | 85.89 | 58.67 | 57.69 | 56.54 | 55.13 | 53.76 | |
| HFDR-MART | 83.43 | 59.73 | 58.96 | 58.73 | 54.56 | 53.42 | ||
| HFDR-AWP | 85.19 | 60.46 | 59.28 | 58.76 | 56.62 | 54.64 | ||
@@
| HFDR-AWP | 59.17 | 36.90 | 36.42 | 36.13 | 32.59 | 31.74 |

### v2

@@ Abstract
Ensuring the robustness of deep neural networks against adversarial attacks remains a fundamental challenge in computer vision. While adversarial training (AT) has emerged as a promising defense strategy, our analysis reveals a critical limitation: AT-trained models exhibit a bias toward low-frequency features while neglecting high-frequency components. This bias is particularly concerning as each frequency component carries distinct and crucial information: low-frequency features encode fundamental structural patterns, while high-frequency features capture intricate details and textures. To address this limitation, we propose High-Frequency Feature Disentanglement and Recalibration (HFDR), a novel module that strategically separates and recalibrates frequency-specific features to capture latent semantic cues. We further introduce frequency attention regularization to harmonize feature extraction across the frequency spectrum and mitigate the inherent low-frequency bias of AT. Extensive experiments demonstrate our method’s superior performance against white-box attacks and transfer attacks, while exhibiting strong generalization capabilities across diverse scenarios.
@@
DNNs have achieved remarkable success in various applications across diverse domains. However, their susceptibility to inconspicuous adversarial perturbations [1, 2, 3] remains a substantial concern. These subtle perturbations, which are undetectable by the human eye, have the potential threat of resulting in erroneous predictions [4, 5]. As a consequence, the security and robustness of DNNs have become areas of widespread concern within both academic and industrial communities. Adversarial training (AT) [6, 7, 8, 9] is widely acknowledged as one of the most effective techniques to enhance adversarial robustness of models. Unlike normal training (NT), which uses clean data to minimize the standard classification loss, AT further incorporates adversarial examples (AEs)—carefully crafted perturbations designed to exploit model vulnerabilities—into the training process.
@@
Despite AT’s remarkable performance, relying solely on AE-based data augmentation proves insufficient as a comprehensive solution. Research has extensively explored regularization [10, 11] and feature denoising [12, 13] methods to enhance adversarial robustness. Recent Fourier analysis-based advances [14, 15, 16, 17] highlight the benefits of incorporating frequency analysis during AT to improve robustness. Building on these insights, Bu et al. [18] developed an adaptive module to control frequency preferences by adjusting low and high-frequency components in feature representations. Similarly, Yucel et al. [19] introduced HybridAugment, a data augmentation method designed to reduce CNNs’ reliance on high-frequency components, followed by HybridAugment++ to unify various frequency-spectrum augmentations. However, the tendency of these approaches to discard or suppress high-frequency components may result in a significant loss of semantic information [15].
We present a comprehensive frequency-domain analysis of adversarial robustness through two experiments illustrated in Figure 1 and Figure 2. Our analysis reveals fundamental differences between normally-trained (NT) and adversarially-trained (AT) models in their utilization of frequency components.
In Figure 1, we analyze the impact of retaining different frequency components on model performance under varying adversarial perturbation strengths. The results reveal distinct behaviors between NT and AT models: (1) For NT models, performance on clean images () improves as more frequency components are included, highlighting the significance of all frequencies in the classification of clean images. However, under adversarial attack conditions (), the model’s performance initially improves with an increase in frequency components, but degrades significantly as more high-frequency features are retained. These results underscore that adversarial attacks primarily affect the high-frequency components, thus affecting the classification accuracy. Notably, the accuracies of the NT model decrease sharply as the strength of adversarial perturbations increases, further emphasizing its sensitivity to attack intensity. (2) For AT models, they exhibit a different characteristic. While their accuracy on clean images is slightly lower than that of NT models, they demonstrate superior robustness under adversarial perturbations. The AT model also experiences performance degradation under adversarial attacks, but the decline is significantly less pronounced than that of the NT model. Furthermore, the classification accuracies of AT models quickly reached saturation, and the inclusion of more high-frequencies has minimal impact on the performance. These findings indicate that the AT model heavily depends on low-frequency components for classification while overlooking high-frequency information. Consequently, adversarial training induces a low-frequency bias.
In Figure 2, we further illustrate the predicted confidence across different frequency components. (1) NT Model: Both high-frequency and low-frequency components from the clean image obtain high post-softmax probabilities. Besides, the original input gets the highest probability, indicating both high-frequency and low-frequency are important for the final probability. For adversarial images, the post-softmax probabilities are significantly dropped across all components, with high-frequency features being more severely impacted, revealing their vulnerability to adversarial perturbations. (2) AT Model: The AT-trained model relies predominantly on low-frequency components for classification in both clean and adversarial images. High-frequency components exhibit notably lower post-softmax probabilities, suggesting that AT-trained models overlook the high-frequency components for the final classification. (3) Comparing the results between NT and AT models, we can find that adversarial training tends to rely on the low-frequency components for the classification, while neglecting the information in the high frequencies of both clean image and adversarial images. This observation is consistent with the low-frequency bias in Figure 1, where the AT models fail to utilize high-frequency features effectively.
Our analysis reveals a fundamental contradiction in how adversarially trained networks process frequency-dependent features. Low-frequency feature robustness manifests as the network’s capacity to effectively capture semantically valuable information for predictions, despite perturbations in the low-frequency domain. Conversely, high-frequency feature robustness reflects the network’s diminished sensitivity to fine-grained details, potentially compromising its ability to capture critical textural nuances [20].
@@
To achieve this, we proposed a novel module “High-Frequency Disentanglement and Recalibration (HFDR)”. Firstly, HFDR employs high-pass filters to generate element-wise frequency attention maps, facilitating the disentanglement of features across different frequency domains. Then through the recalibration of vulnerable high-frequency features, the model effectively captures potential semantic cues within the high-frequency domain. The fusion of recalibrated high-frequency features with low-frequency features enhances the network’s ability to extract distinct frequency features. To further mitigate the low-frequency biases induced by AT, we introduce a frequency domain attention regularization. This regularization strategy aims to harmonize the network’s extraction capacity for different frequency domain information. Notably, our proposed model achieves comparable or superior robustness compared to state-of-the-art methods, while incurring minimal additional computational overhead. Addtionally, the results of our model in Figure 2 show a significant increase in post-softmax probabilities associated with the high-frequency components.
@@
This study identifies a significant low-frequency bias in conventional adversarial training methods, which often overlook high-frequency features containing valuable semantic information.
We propose a High-Frequency Disentanglement and Recalibration (HFDR) module that extracts latent high-frequency cues, integrates frequency attention mechanisms, and addresses low-frequency bias to improve model robustness.
The proposed model achieves state-of-the-art robustness with minimal additional computational cost and can be seamlessly integrated with existing adversarial training methods to further enhance model performance.
@@ II-A Adversarial Attack
Deep neural networks (DNNs) demonstrate remarkable performance in various tasks by learning intricate relationships between inputs and outputs through complex, non-linear, and high-dimensional mappings. However, DNNs are susceptible to adversarial attacks, in which carefully crafted perturbation can deceive the model, leading to incorrect predictions. Several research studies have been dedicated to investigating the vulnerability of models and have proposed various advanced adversarial attack methods. Cao et al. [21] has demonstrated that adversarial attacks pose substantial real-world threats to critical applications including authentication systems and physiological monitoring services.
@@ II-B Adversarial Training Defense Methods
The objective of the inner maximization is to generate adversarial examples, including attacks such as the PGD attack [22] and other attacks [27, 28, 29, 30]. The outer minimization aims to develop effective training strategies for optimizing network parameters, thereby enhancing the model’s robustness against attacks. AdvProp [31] enhances robustness by using adversarial examples with separate batch normalization to address distribution differences, improving generalization in image recognition tasks. Universal adversarial training [32] enhances robustness by incorporating universal adversarial perturbations during training, modeling the problem as a two-player min-max game. Rice et al. [33] conducted a study on the occurrence of overfitting in robust adversarial training and proposed employing a validation set protocol while performing model selection (PGD-AT):
Zhang et al. [34] categorized robust errors into “natural errors” and “boundary errors”, balancing robustness and accuracy by controlling the Kullback–Leibler divergence between clean and adversarial outputs (TRADES):
Wang et al. [35] introduced loss weighting technical based on the classification status to mitigate adversarial misclassification errors while maintaining robustness (MART):
Wu et al. [36] proposed the flattening adversarial weight loss landscape, enhancing model robustness by directly perturbing network parameters and addressing adversarial vulnerabilities in both input and parameter spaces (AWP):
Jia et al. [6] introduced a computationally intensive method that involves automatically generating attacks from a strategy network and strategy set to adaptively find optimal adversarial strategies (LAS-AT):
These adversarial training approaches exhibit distinct characteristics in their robustness optimization strategies. TRADES and MART represent complementary advances in leveraging clean image properties: TRADES optimizes the KL divergence to balance accuracy and robustness, while MART enhances this framework through adaptive confidence-based weighting for misclassified samples. AWP introduces a distinct paradigm by addressing vulnerability in the parameter space via weight perturbations, moving beyond traditional input-space approaches. LAS-AT further advances the field by introducing a learnable attack strategy network that dynamically adapts adversarial examples during training, representing a significant shift toward adaptive defense mechanisms. While these approaches have advanced the field through novel attack strategies and training optimizations, they primarily focus on the training process itself, leaving the crucial aspect of feature representation properties during network inference relatively unexplored.
@@ III Methodology
Adversarial training can introduce an inherent low-frequency bias, limiting the model’s ability to capture meaningful semantic information across different frequency domains. This bias arises from the concentration of adversarial perturbations primarily in the high-frequency domain. To address this issue, we propose a novel High-Frequency Disentanglement and Recalibration (HFDR) module, as illustrated in Figure 3. The HFDR module consists of three key components: feature disentanglement, high-frequency feature recalibration, and frequency-based attention regularization.
In the feature disentanglement phase (Section III-A), we generate feature attention maps that selectively focus on distinct frequency domains, enabling the separation of features into high-frequency and low-frequency components. The high-frequency feature recalibration stage (Section III-B) then refines the high-frequency information to effectively capture valuable semantic features. Additionally, we introduce a novel regularization technique, Frequency Attention Regularization, to mitigate the low-frequency bias induced by adversarial training (Section III-C).
@@ III-A Feature Disentanglement
Feature disentanglement involves separating and isolating distinct components within a feature representation at a specific layer. In network inference, let denote the feature maps at a predefined layer, where , , and represent the channel, height, and width dimensions, respectively. Recent approaches commonly rely on low-pass or high-pass filters for feature disentanglement; however, these methods often result in irreversible loss of information [39, 40]. To address this limitation, we employ an SRM filter [41] to extract frequency-specific features while ensuring that the proposed disentanglement process retains both high- and low-frequency components:
where represents the high-frequency feature components aligned across channels through a convolution, and represents the low-frequency components calculated as the residual of the original feature maps after subtracting the high-frequency components.
To achieve feature disentanglement while minimizing information loss, we propose using filters to generate element-wise attention maps for both high-frequency (HF) and low-frequency (LF) semantic sub-features. These attention maps are then used to disentangle the features. To address the potential problem that the enhanced model robustness is due to gradient masking [42], we incorporate the Gumbel-Softmax method [43] to generate differentiable attention maps specifically for high-frequency components:
where represents distinct semantic sub-features, and denotes the sigmoid operation. The term incorporates Gumbel noise, defined as with . We set to a fixed value for all samples to control the sampling randomness. The term governs the impact of . Additionally, the attention maps for low-frequency features can be computed as .
@@ III-C Frequency Attention Regularization
Feature maps inherently contain multiple frequency components: low-frequency features represent global structures, while high-frequency components capture fine-grained details. In adversarial training, however, prior methods [33, 36, 35, 18, 48] often induce a low-frequency bias, leading the model to favor low-frequency features and overlook high-frequency information. This bias limits the model’s ability to robustly utilize both types of features.
To address this fundamental limitation, we propose the Frequency Attention Regularization (FAR) mechanism, a novel approach that dynamically calibrates the model’s attention across the frequency spectrum. FAR introduces a principled framework for maintaining an optimal balance between high- and low-frequency feature extraction, effectively counteracting the inherent low-frequency bias in adversarial training. Our empirical analysis demonstrates that this balanced feature utilization significantly enhances the model’s robustness against diverse adversarial attacks.
For each channel, we compute a high-frequency attention map, defined as . The FAR regularization term is given by:
where denotes the high-frequency attention map for the -th channel of the -th image, represents the batch size, and indicates the number of channels in the feature maps processed by the HFDR module. The parameter serves as a critical balance factor that adaptively controls the ratio between high- and low-frequency feature attention, while determines the norm used in penalizing deviations from the target ratio.
The FAR mechanism operates through two key components: (1) the adaptive balance factor , which precisely modulates the emphasis on high-frequency features during network inference, preventing scenarios where fails to adequately capture important high-frequency patterns, and (2) the p-norm penalty, which governs the intensity of correction applied when the ratio between high-frequency and low-frequency features deviates from the target distribution.
@@ III-D Model Training
where controls the influence of , and represents the loss function associated with various AT models.
By incorporating the HFDR module, we enhance the model’s capacity to extract and utilize features across the entire frequency spectrum, leading to improved adversarial robustness. The complete training procedure is detailed in Algorithm Algorithm 1, which provides a step-by-step description of our approach, including the initialization of the HFDR module, the generation of adversarial examples, and the optimization of the combined loss function.
@@ IV-A Experimental Setup
| SGLR [51] | CVPR-24 | 85.81 | 57.83 | 57.40 | 56.99 | 54.63 | 52.70 | |
| HFDR-AT | Ours | 85.89 | 58.67 | 57.69 | 56.94 | 55.13 | 53.76 | |
| HFDR-TRADES | 84.18 | 58.99 | 58.40 | 58.31 | 55.43 | 53.86 | ||
| HFDR-MART | 83.43 | 59.73 | 58.96 | 58.73 | 54.56 | 53.92 | ||
| HFDR-AT-AWP | 85.19 | 60.46 | 59.28 | 58.76 | 56.62 | 54.64 | ||
@@
| HFDR-AT-AWP | 60.97 | 36.95 | 36.47 | 36.16 | 32.70 | 31.54 |
Implementation Details: We evaluate the robustness of our approach by conducting experiments on the CIFAR-10 [52], CIFAR-100 [52], and Imagenette [53] datasets, utilizing the WideResNet34-10 (WRN34-10) [54], AlexNet [17], and ResNet-18 (RN-18) [55] as the baseline models. For the implementation, we insert the HFDR module after the initial convolutional layer. To evaluate the impact of HFDR on improving the robustness of various AT techniques, our model was combined with three AT methods (PGD-AT [33], MART [35], AWP [36]), denoted as, HFDR-AT, HFDR-MART, and HFDR-AT-AWP. We compared our approach with several baseline methods, including, PGD-AT [33], TRADES [34], AWP [36], MART [35], LAS-AT [6], FSR [49], FPCM [18], CFA [50] and SGLR [51].
We trained WRN34-10 with identical hyperparameters and training details as specified in the original paper [33, 35, 36]. To ensure a fair comparison, we adopted the configuration of PGD-AT [33] as used in the experimental settings of FPCM [18], FSR [49] and CFA [50]. For ResNet18 and AlexNet architectures, we initialized the learning rate to 0.1. The learning rate was subsequently reduced by a factor of 0.1 at the 90th and 95th epochs. The SGD optimizer with a momentum of 0.9 and a weight decay factor of 5e-4 was employed for optimization. In the HFDR hyperparameter settings, we set and for the loss term in Equation 11. The experiments were conducted on two NVIDIA RTX-A4000 GPUs.